Alpha Company, on March 1, 2017 has a beginning Work in Process inventory of zero. All materials are added into production at the beginning of its production. There is only one production WIP inventory. During the month 100,000 units were started. At the end of the month all started units were 80% complete with respect to conversion. Direct Materials placed into production had a total cost of $24,000 and the total conversion cost for the month was $16,000. Alpha uses the weighted-average process costing method. Use this information to determine for the month of March: (Round & enter final answers to the nearest penny.)

1. the cost per equivalent unit of direct material

2. the cost per equivalent unit of conversion
Answer #1

Calculation of equivalent units

Since material is placed at the beginning, hence it will be equal to 100%. So equivalent units for material will be completed units + closing WIP

= 100000 units

Conversion cost

= (Units completed*100%) + (closing WIP*80%)

= (0*100%) + (100000*80%)

= 80000 units (because no unit is completed)

Equivalent cost per unit

= Cost involved/ equivalent units

Material = $24000/100000 units = $0.24 per unit

Conversion cost = $16000/80000 units = $0.20 per unit

Feel free to ask any queries..

Also plz upvote it means a lot.... Thank you